If A and B are independent events with P(A) = .2 and P(B) = .6, then P(AB) = _____.

a. .12
b. .68
c. .62
d. .60
Answer #1

Given that PCA)= 0.2 PCB) = 0.6 first calculate PCA NB) = P(A) XP (B) = 6.2 2016 = 0.12 Now, PCA UB) = P(A) + PCB) - P(ANB) =